7. At night a watchman shines a torch at a pool of water and observes a reflected spot of light on the opposite wall. The angle of depression of the torch and ray is 30o, and P, the point of reflection from the water surface, is 6 m from the wall.Calculate the distance h of the spot of light above the water

8. You are flying in a light aircraft at a speed of 39 ms-1. An eagle, whose speed is 18 ms-1, is flying directly toward you. Each of the given speeds is relative to the ground. The eagle emits a shrill cry whose frequency is 3400 Hz. What frequency do you hear? (Assume Speed of sound is 345 ms-1)
9. A man enters a tall tower and notes that a long pendulum extends from the ceiling to half the height of the building. If the height of the building is 60 m determine the period of oscillation?
1
Expert's answer
2019-09-30T10:09:11-0400

7.


"h=d\tan{30\degree}"

"h=6\tan{30\degree}=3.5\ m"

8.


"f'=f\frac{v+v_o}{v-v_s}"

"f'=3400\frac{345+39}{345-18}=4000\ Hz"

9.


"T=2\pi \sqrt{\frac{0.5l}{g}}"


"T=\pi \sqrt{\frac{2l}{g}}=\pi \sqrt{\frac{2(60)}{9.8}}=11\ s"